# Equipping Sketch Patches with Context-Aware Positional Encoding for Graphic Sketch Representation

### v1

@@ 5 Conclusions
We have presented DC-gra2seq to encode sketch drawing orders by PE for learning context-aware graphic sketch representations. Each sketch patch is equipped with a sinusoidal absolute PE, indicating where it is drawn in a drawing order, and a learnable relative PE, restoring the sequential information from the contextual positions of sketch components. The attendance of PEs encourages DC-gra2seq to simultaneously consider both the visual patterns from the canvas and the sequential positions from the drawing order, which benefits sketch representation learning verified on controllable sketch synthesis and sketch healing.
@@ 6 Acknowledgments
This work was supported by the Joint Funds of the National Natural Science Foundation of China (U2033218).

### v2

@@ 5 Conclusions
We have presented DC-gra2seq to encode sketch drawing orders by PE for learning context-aware graphic sketch representations. Each sketch patch is equipped with a sinusoidal absolute PE, indicating when it is drawn in a drawing order, and a learnable relative PE, encoding the contextual relationships between sketch patches. Both types of PEs are injected into graph nodes instead of edges to learn robust sketch representation against variants of sketch drawings. The attendance of PEs encourages DC-gra2seq to simultaneously consider the visual patterns from the canvas, the sequential positions from drawing orders and the unseen contextual information between patches, which benefits sketch representation learning verified on controllable sketch synthesis and sketch healing.
@@ 6 Acknowledgments
This work was supported by the Fundamental Research Funds for the Central Universities (2232024D-28) and the National Natural Science Foundation of China (62406064).